Find the Call Price and then the Put Price Using Black-Scholes model CBarede FINANCO dlchPrice 50 Eseaver roe 450 heh ree Aate 5% Tue Natuidye lentanke laaded desiatona0% Rak Prce Cell Rice UBWR
Answer #1

Civen data is S = 45 K= 45 ö 2=31307 4h R=45 rrios T-15 (6/12] 8=:05 T: 05 (o values in this The price of puta Call option :value of put options P = Exert n(d) - SON(-d) . - 6x ē XN6.01) - 45XNV--22) 468.94560 x.496 - 454.4189. = 27.762 – 18.5806 =